OpenStudy (anonymous):

The area (A) of a ring is A = π r2 - π s2. Find the area A if π = 3.14, r = 6 and s = 4. A = 62.8 A = 197.192 A = 31.4 A = 6.28

OpenStudy (unklerhaukus):

\[A=3.14\times6^2-3.14\times4^2\]

OpenStudy (unklerhaukus):

\[A=3.14(6^2−4^2)\]

OpenStudy (anonymous):

A = π r2 - π s2 A=(3.14)(6^2)-(3.14)(4^2) A=113.04-50.24 A=62.8 so option 1 is the answer